# Off-Policy Evaluation for Sequential Persuasion Process with Unobserved Confounding

### v1

@@ IV Construction of POMDP Structure
In this POMDP, the unobserved state $x_{t}$ plays the role of an unobserved confounder between the action $u_{t}$ and the next state $x_{t+1}$. Since the exact dynamics of the POMDP are unknown and an unobserved confounder exists, we do not have access to $p^{\bm{g}}_{i}(\cdot)$ or $p^{\bm{g},o}_{t}(\cdot)$. Consequently, we can only hope to compute $J(\bm{g})$ based on a logged dataset. To this end, we consider that we have access to a dataset $\bm{D}^{b}=\{\tau^{o,k}_{T+1}\}_{k=0}^{n}$, which is a collection of n-observable trajectories, generated by an unknown behavioral control strategy $\bm{g}^{b}$. Typically, OPE techniques are used to compute the performance of any given strategy based on $\bm{D}^{b}$. However, the unobserved confounder, by its nature, poses a challenge to OPE. Furthermore, we demonstrate how the performance of any given evaluation strategy $\bm{g}^{e}$ be computed based on $\mathcal{D}^{b}$ and how the issue of unobserved confounding is adjusted by using the proximal learning technique.
@@ V-B Evaluation using proximal learning
In the following analysis, we demonstrate the proximal learning method for sequential decision-making problems presented in [15] for the reformulated POMDP. The essence of proximal learning is to find proxies to infer the average influence of the unobserved confounder on the performance of a strategy. Firstly, we present a result that provides a method for computing the distribution $p^{\bm{e}}(\cdot)$ on the reward $r$ at each time $t$ as a function of just the observable trajectory $\tau^{o}_{t}$. In the proof, we show how past and current observations of the state can be used as proxies at each $t$, to infer the performance of an evaluation strategy. For the exposition, we consider the following assumptions.
@@
At each $t=1,\ldots,T$, the conditional matrices $P^{\bm{b}}(Y_{t}\mid Y_{t-1},u_{t})$ and $P^{\bm{b}}(X_{t}\mid Y_{t-1},u_{t})$ are invertible.
@@
The proof proceeds in two steps. We expand the expression for the reward distribution and decompose it into two components based on dependence on the strategy. The first component includes strategy-dependent distributions on observed variables, which can be computed given a specific evaluation strategy $\bm{g}^{e}$. On the other hand, strategy-independent distributions can be over a combination of observable and unobservable variables. Next, we show how to compute the strategy-independent component from observational data. This is achieved by using proxy variables to construct the weight matrices $W_{t}(\tau^{o}_{t})$ for any observable trajectory $\tau^{o}_{t}$.
We begin by expanding the strategy dependent distribution $p^{\bm{e}}(\tau_{t+1})$ for any $\tau_{t}=(u_{0:t},y_{0:t},x_{0:1})$ in (14) as
@@
where the right-hand side of this equation is the weight matrix $W_{k}(\tau^{o}_{k})$. This shows that the consecutive multiplication of terms at indexes $k$ and $k+1$ is equivalent to the multiplication of weight matrices. Substituting the weight matrix in (38) into the expression for reward distribution in (26) will give us the result presented in Theorem 2. ∎
VI Application Example
In this section, we elaborate on a framework based on SPP to elucidate how AI can communicate effectively within the context of industrial safety management. This framework aims to influence the behavior of warehouse delivery associates engaged in the transportation of packages through various means, including forklifts, conveyor systems, and manual handling. Packages may carry specific handling instructions, such as “fragile,” “temperature-sensitive,” or “do not turn upside down.” The primary objective for associates is to maximize operational efficiency to enhance their incentives, while safety managers must ensure both package throughput and the safety of workers. The daily demand and the inherent characteristics of the packages define the system’s current state. Equipped with advanced computational resources, safety managers can accurately estimate the distribution from which this state is derived, thereby possessing comprehensive knowledge of the state distribution—information not directly accessible to the associates. To facilitate safe practices, safety managers communicate with associates via personal devices, dispatching signals that range from “extremely cautious” to “at ease.” These signals are critical as they guide associates in updating their beliefs before decision-making. For instance, on high-demand days, cautionary signals are prevalent, whereas “at ease” messages are more common during periods of lower demand. It is essential to recognize that individual risk tolerance influences how associates interpret these signals; a risk-averse associate may equate a cautious signal with a high-demand scenario, while a risk-seeking associate might undervalue caution even in similar contexts. This variability underscores the importance of considering past interactions and rewards, which can significantly shape the interpretation and effectiveness of safety communication strategies.
VII Conclusion
In this paper, we developed a framework to study an SPP that incorporates unobserved confounding affecting the belief updates of the receiver. We formulated an equivalent POMDP model, where the optimal observation-based control strategy corresponds to the optimal meta-policy for the SPP. Furthermore, we demonstrated how to conduct off-policy evaluation of sender-designed signaling strategies, illustrating that the use of proxy variables in the equivalent model allows for performance evaluation in the presence of unobserved confounding. We provided an example involving interactions between an industrial safety manager and an in-house delivery associate to illustrate a practical application of our framework.
While this work contributes valuable insights, several limitations warrant attention in future research. These include enhancing computational tractability, integrating OPE techniques such as sensitivity analysis [18] and bridge functions [19], extending to continuous policy and belief spaces, optimizing for meta-policies [20], and applying the framework to complex human decision-making tasks.
Acknowledgments
The authors would like to thank Dr. Aditya Dave for valuable discussions on the formulation and solution approach of this work.

### v2

@@ IV Construction of POMDP Structure
In this POMDP, the unobserved state $x_{t}$ plays the role of an unobserved confounder between the action $u_{t}$ and the next state $x_{t+1}$. Since the exact dynamics of the POMDP are unknown and an unobserved confounder exists, we do not have access to $p^{\bm{g}}_{i}(\cdot)$ or $p^{\bm{g},o}_{t}(\cdot)$. Consequently, we can only hope to compute $J(\bm{g})$ based on a logged dataset. To this end, we consider that we have access to a dataset $\bm{D}^{b}=\{\tau^{o,k}_{T+1}\}_{k=0}^{n}$, which is a collection of n-observable trajectories, generated by an unknown behavioral control strategy $\bm{g}^{b}$. Typically, OPE techniques are used to compute the performance of any given strategy based on $\bm{D}^{b}$. However, the unobserved confounder, by its nature, poses a challenge to OPE. In the next section, we analyze how unobserved confounding in this context is handled.
@@ V-B Evaluation using proximal learning
In the following analysis, we demonstrate the proximal learning method for sequential decision-making problems presented in [15] for the reformulated POMDP. For the exposition, we consider the following assumptions.
@@
At each $t=1,\ldots,T$, the matrices $P^{\bm{b}}(Y_{t}\mid Y_{t-1},u_{t})$ and $P^{\bm{b}}(X_{t}\mid Y_{t-1},u_{t})$ are invertible.
@@
The proof proceeds in two steps. We expand the expression for the reward distribution and decompose it into two components based on dependence on the strategy. The first component includes strategy-dependent distributions on observed variables, which can be computed given a specific evaluation strategy $\bm{g}^{e}$. On the other hand, strategy-independent distributions can be over a combination of observable and unobservable variables. Next, we show how to compute the strategy-independent component from observational data using proximal learning. Details of the mathematical arguments are provided in Appendix A. ∎
VI Application Example
In this section, we elaborate on a framework based on SPP to elucidate how AI can communicate effectively within the context of industrial safety management. This framework aims to influence the behavior of warehouse associates engaged in the transportation of packages through various means, including forklifts, conveyor systems, and manual handling. Packages may carry specific handling instructions, such as “fragile,” “temperature-sensitive,” or “do not turn upside down.” The primary objective for associates is to maximize operational efficiency to enhance their incentives, while safety managers must ensure both package throughput and the safety of workers. The daily demand and the inherent characteristics of the packages define the system’s current state. To facilitate safe practices, safety managers communicate with associates via personal devices, dispatching signals that range from “extremely cautious” to “at ease.” These signals are critical as they guide associates in updating their beliefs before decision-making. For instance, on high-demand days, cautionary signals are prevalent, whereas “at ease” messages are more common during periods of lower demand. It is essential to recognize that individual risk tolerance influences how associates interpret these signals; a risk-averse associate may equate a cautious signal with a high-demand scenario, while a risk-seeking associate might undervalue caution even in similar contexts. This variability underscores the importance of considering past interactions and rewards, which can significantly shape the interpretation and effectiveness of safety communication strategies.
VII Conclusion
In this paper, we developed a framework to study an SPP that incorporates unobserved confounding affecting the belief updates of the receiver. We formulated an equivalent POMDP model, where the optimal observation-based control strategy corresponds to the optimal meta-policy for the SPP. Furthermore, we demonstrate off-policy evaluation using proximal learning to showcase policy evaluation under unobserved confounding. We provide an industrial safety- management example to illustrate practical applications.
Future works could consider enhancing computational tractability for implementation, integrating different OPE techniques (sensitivity analysis [18], bridge functions [19]), and extend this framework to continuous spaces to be applicable to complex human decision-making tasks.
Acknowledgments
The authors would like to thank Dr. Aditya Dave for valuable discussions on the formulation and solution approach.
Appendix A - Proof of Theorem 2
The essence of proximal learning is to find proxies to infer the average influence of the unobserved confounder on the performance of a strategy. Firstly, we present a result that provides a method for computing the distribution $p^{\bm{e}}(\cdot)$ on the reward $r$ at each time $t$ as a function of just the observable trajectory $\tau^{o}_{t}$. We present a detailed proof of Theorem 2 and show how past and current observations of the state can be used as proxies at each $t$, to infer the performance of an evaluation strategy.
We begin by expanding the strategy dependent distribution $p^{\bm{e}}(\tau_{t+1})$ for any $\tau_{t}=(u_{0:t},y_{0:t},x_{0:t})$ in (6) as
@@ Acknowledgments
where the right-hand side of this equation is the weight matrix $W_{k}(\tau^{o}_{k})$. This shows that the consecutive multiplication of terms at indexes $k$ and $k+1$ is equivalent to the multiplication of weight matrices. Substituting the weight matrix in (30) into the expression for reward distribution in (18) will give us the result presented in Theorem 2.
Appendix B - Discussions
Finite Belief Space: In the context of human-AI recommendation systems, the AI communicates with a human who serves as the receiver. Typically, humans interpret and process information from these recommendations or signals in categorical terms rather than as nuanced probability distributions. For instance, an individual’s belief about the likelihood of rain on a given day may range from low to medium to high, reflecting their level of confidence. Additionally, due to computational constraints, a receiver may simplify their belief space by utilizing a finite set of categories, thus reducing the complexity involved in updating their beliefs to make self-interested decisions.
Information vector: In this work, we have considered that the receiver’s reward realizations are observable by the sender. Consequently, the information vector includes $(\rho^{r}_{0:i-1})$. Based on this definition, we define the corresponding notion of state for the POMDP. However, in settings where the receiver’s reward need not be observable to the sender, the information vector must be redefined to exclude this information. In such cases, the notion of state can be adapted accordingly, such that the Markovian dynamics of the system is preserved.
Invertibility of statistically computed matrices:     The assumption that the matrices $P^{\bm{b}}(Y_{t}\mid Y_{t-1},u_{t})$ and $P^{\bm{b}}(X_{t}\mid Y_{t-1},u_{t})$ are invertible may not hold when they are computed from data. The quality of available offline data plays a crucial role to determine the validity of this assumption. However, as shown in [15], the invertibility of these matrices holds without explicitly accounting for their empirical estimation. This motivates the consideration of a set of matrices that are both invertible and sufficiently close to the empirical estimates. In this case, a robust estimate of off-policy evaluation could be useful, which can serve as an avenue for future research.